OpenStudy (anonymous):

Find the 100th term of the sequence -7, 3, 13, 23, ... -1097 1093 40

OpenStudy (luigi0210):

Do you know the arithmetic sequence formula?

OpenStudy (anonymous):

Oh and 1083

OpenStudy (anonymous):

Yes! If a sequence of values follows a pattern of adding a fixed amount from one term to the next. (I looked up earlier)

OpenStudy (luigi0210):

\[\LARGE a_{n}=a_{1}+(n-1)d\] where a1 is the first term of the sequence, d is the common difference, n is the number of the term to find.

OpenStudy (anonymous):

Yup that's it!

OpenStudy (luigi0210):

So now just plug in your information.. \[\LARGE a_{100}=-7+(100-1)10\]

OpenStudy (luigi0210):

So solve this: \[\LARGE a_{100}=-7+990\]

OpenStudy (luigi0210):

I messed up didn't I? =.=

OpenStudy (anonymous):

893

OpenStudy (anonymous):

Lol yeaa it's okay

OpenStudy (anonymous):

So what am I plugging in? Tell me that and I'll be good to go! :)

OpenStudy (luigi0210):

I did everything right tho o.O n=100 a1=-7 d=10

OpenStudy (anonymous):

Hmm, Let's see. Where did 10 come from?

OpenStudy (luigi0210):

That's the commone difference, is it not? It's being add everything time.

OpenStudy (anonymous):

Wow Correct, and my answer choices are: 1083 -1097 1093 40 Something's wrong with my choices aren't there?

OpenStudy (anonymous):

*facepalm* Of course it is!

OpenStudy (anonymous):

It's not "Find the 100th term of the sequence" it's "Find the 110th term of the sequence!"

OpenStudy (anonymous):

So that means it will be a110=-7+(110-1)10 ! Right?

OpenStudy (luigi0210):

Oh, that would make more sense xD But yea, that's right.. it should give you 1083 :)

OpenStudy (anonymous):

And then 1083!

OpenStudy (anonymous):

Right! Thank you Luigi!

OpenStudy (luigi0210):

You're welcome~

OpenStudy (anonymous):

See ya!